Krein’s differential system and its generalization

Lev Sakhnovich Courant Institute, USA

In the famous M.G. Krein’s paper( Dokl.Akad. Nauk, 105,637-640,1955) a special important class of differential systems was considered.M.G. Krein announced a number of fundamental facts of the direct and inverse spectral theory of this class.Unfortunately these important results where published without proof.We proved an essential part of assertions stated in this paper and generalized them to a broad class of canonical differential systems.